# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v2

@@ Appendix G Experiment Results of Detecting 2024 Olympic News or Machine-generated News
Figure 7: Comparisons between our method and baselines for detecting 2024 Olympic news and machine-generated news. Fake news are generated by 3 source models: Gemini-1.5-Flash, Gemini-1.5-Pro and PaLM 2. The scoring model used is Neo-2.7 with the score function of Fast-DetectGPT. We consider five batch sizes: $k=25,50,100,250,500$. The left column displays results from the permutation test without correction, while the right column presents results of permutation test with corrected significance levels $\alpha$ for each batch test.
@@
Appendix H Experiment Results of detecting texts from three domains
@@
We also test on the dataset of Bao et al. [2023] to explore the influence of text domains on the detection result of our algorithm. In this experiment, we only consider Scenario 1. We let $x_{t}$ be human-written text from three datasets following Mitchell et al. [2023], each chosen to represent a typical LLMs application scenario. Specifically, we incorporate news articles sourced from the XSum dataset [Narayan et al., 2018], stories from Reddit WritingPrompts dataset [Fan et al., 2018] and long-form answers written by human experts from the PubMedQA dataset [Jin et al., 2019]. Then, the capability of our algorithm is evaluated by detecting the source of texts $y_{t}$ originated from the above source models or human datasets. Source models involved in this experiment are GPT-3  [Brown, 2020], ChatGPT  [OpenAI, 2022], and GPT-4  [Achiam et al., 2023] while the scoring model is Neo-2.7 [Black et al., 2021]. The perturbation function for DetectGPT and NPR is T5-11B, and the sampling model for Fast-DetectGPT is GPT-J-6B.
@@
Figure 8(a) and Figure 8(b) show the averaged results when two streams of texts are from the same domain and different domains respectively, i.e., both sequences of texts are sampled from/prompted according to the same dataset/different datasets. Specifically, we get the results by averaging rejection time and FPR under each $\alpha$-significance level across the corresponding results of three LLMs (GPT-3, ChatGPT, GPT-4) and three same/different domain settings shown in Figure 10 and 8(b). After comparing these two figures, we find that the correct declaration of an LLM source happens more quickly when the prepared texts and the texts to be detected are from the same domain. Our algorithm can control FPRs below the significance level $\alpha$ for all score functions while the averaged rejection times for them are all shorter than 150. Among all functions, Fast-DetectGPT is the fastest to reject $H_{0}$, with average value of $\tau$ being around 40 for same-domain texts and about 80 for different-domain texts.
@@
(a) Same domain: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from XSum or generated by GPT-3 based on XSum.
(b) Same domain: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from XSum or generated by ChatGPT based on XSum.
(c) Same domain: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from XSum or generated by GPT-4 based on XSum.
(d) Same domain: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from Writing or generated by GPT-3 based on Writing.
(e) Same domain: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from Writing or generated by ChatGPT based on Writing.
(f) Same domain: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from Writing or generated by GPT-4 based on Writing.
(g) Same domain: $x_{t}$ is sampled from PubMed, and $y_{t}$ is sampled from PubMed or generated by GPT-3 based on PubMed.
(h) Same domain: $x_{t}$ is sampled from PubMed, and $y_{t}$ is sampled from PubMed or generated by ChatGPT based on PubMed.
(i) Same domain: $x_{t}$ is sampled from PubMed, and $y_{t}$ is sampled from PubMed or generated by GPT-4 based on PubMed.
@@
Figure 9: Test results: mean rejection times (under $H_{1}$) and FPRs (under $H_{0}$) of each significance level $\alpha$ for 10 score functions, with texts $x_{t}$ and $y_{t}$ from the same domain. There are 3 source models: GPT-3, ChatGPT and GPT-4.
(a) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from Writing or generated by GPT-3 based on Writing.
(b) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from Writing or generated by ChatGPT based on Writing.
(c) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from Writing or generated by GPT-4 based on Writing.
(d) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from PubMed or generated by GPT-3 based on PubMed.
(e) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from PubMed or generated by ChatGPT based on PubMed.
(f) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from PubMed or generated by GPT-4 based on PubMed.
(g) Different domains: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from PubMed or generated by GPT-3 based on PubMed.
(h) Different domains: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from PubMed or generated by ChatGPT based on PubMed.
(i) Different domains: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from PubMed or generated by GPT-4 based on PubMed.
@@
Figure 10: Test results: mean rejection times (under $H_{1}$) and FPRs (under $H_{0}$) of each significance level $\alpha$ for 10 score functions, with texts $x_{t}$ and $y_{t}$ from different domains. There are 3 source models: GPT-3, ChatGPT and GPT-4.
@@
Since we let $\epsilon$ equal to the actual $\Delta$ value of two sequences of human texts, which ensures FPRs of all score functions for each significance level $\alpha$ remain below $\alpha$. As we have mentioned previously, our algorithm can ensure a nonnegative supermartingale wealth under $H_{0}$ and thus control the type-I error. Besides, the relative magnitude of $(\Delta-\epsilon)$ and $(d_{t}-\epsilon)$ would influence the rejection time under $H_{1}$.
@@
Table 9: Values of $d_{t}$ with the assumption that we know the range of $g_{t}=\lvert\phi(x_{t})-\phi(y_{t})\rvert$ beforehand. Every two columns starting from the third column represent the $d_{t}$ values for each $t$ used under $H_{1}$ and $H_{0}$ in each test scenario. For instance, the values in the third column are calculated as $\max_{i,j\leq 150}\lvert\phi(x_{i})-\phi(y_{j})\rvert$, where $\phi(x_{i})$ is the score of $i$-th XSum text and $\phi(y_{j})$ is the score of the $j$-th text generated by GPT-3 based on XSum. Similarly, the fourth column calculates the maximum difference between the scores of two text sequences which are both sampled from XSum dataset. The $d_{t}$ values is then used to define the domain of $\theta_{t}$ in our algorithm, where $\theta_{t}\in[-1/2d_{t},0]$.
@@
Table 10: Values of $d_{t}$ with the assumption that we know the range of $g_{t}=\lvert\phi(x_{t})-\phi(y_{t})\rvert$ beforehand. Every two columns starting from the third column represent the $d_{t}$ values for each $t$ used under $H_{1}$ and $H_{0}$ in each test scenario. For instance, the values in the third column are calculated as $\max_{i,j\leq 150}\lvert\phi(x_{i})-\phi(y_{j})\rvert$, where $\phi(x_{i})$ is the score of $i$-th XSum text and $\phi(y_{j})$ is the score of the $j$-th text generated by GPT-3 based on Writing. Similarly, the fourth column calculates the maximum difference between the scores of two text sequences with $x_{i}$ sampled from XSum and $y_{j}$ sampled from Writing dataset. The $d_{t}$ values is then used to define the domain of $\theta_{t}$ in our algorithm, where $\theta_{t}\in[-1/2d_{t},0]$.
@@
Table 11: Values of $\Delta$ derived by using Neo-2.7 as the scoring model for the first eight score functions to score texts from the same domain. There are three source models: GPT-3, ChatGPT and GPT-4. Every two columns starting from the third column represent the $\Delta$ values under $H_{1}$ and $H_{0}$ in each test scenario. For instance, the third column presents the average absolute difference in scores between 150 Xsum texts and 150 texts generated by GPT-3 based on XSum texts. The fourth column illustrates the $\Delta$ value between 150 XSum texts and 150 XSum texts.
@@
Table 12: Values of $\Delta$ derived by using Neo-2.7 as the scoring model for the first eight score functions to score texts from different domains. There are three source models: GPT-3, ChatGPT and GPT-4. Every two columns starting from the third column represent the $\Delta$ values under $H_{1}$ and $H_{0}$ in each test scenario. For instance, the third column presents the average absolute difference in scores between 150 Xsum texts and 150 texts generated by GPT-3 based on Writing texts. The fourth column illustrates the $\Delta$ value between 150 XSum texts and 150 XSum texts.
@@ Appendix I Addressing Parameter Estimation Challenges and Mixed LLMs or Sources Tasks
Estimate Parameters Based on More Samples. Better parameter estimation can enhance the performance of our algorithm. For instance, in the previous experiments, we used the first $10$ samples to estimate parameters $d_{t}$ and $\epsilon$. Here, we give an example to show that using a larger sample size for estimation could possibly yield better results. Specifically, using $20$ samples for estimation, with test begins at the $21$-th time step, could lead to improved algorithm performance.
@@
Figure 11: Comparison of different durations in the initial stage for parameter estimation in Scenario 2. Here, text $x_{t}$ is sampled from XSum and $y_{t}$ is from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Neo-2.7. Better parameter estimation can enhance the performance of our algorithm. In the previous experiments, we used the first $10$ samples to estimate parameters $d_{t}$ and $\epsilon$, as shown in (a). The subfigure (b) suggests that a longer duration of for parameter estimation could possibly yield better results, where we emphasize that the test begins at $t=21$ after estimating the parameters.
@@
Figure 12: (Extension to other settings) (a) Results when the sequence of texts $y_{t}$ are produced by various LLMs instead of a single one. (b) Results under the setting that the null hypothesis corresponds to the case that all the texts from the unknown source are human-written, while the alternative hypothesis corresponds to the one that not all $y_{t}$ are human-written.
The above results reflect a real-world scenario where parameter values are estimated from the first $10$ samples. The performance of our method could be further enhanced with prior knowledge of the parameters.

### v3

@@ Appendix G Experimental Results of Detecting 2024 Olympic News or Machine-generated News
Figure 7: Comparisons between our method and two baselines for detecting 2024 Olympic news and machine-generated news. Fake news are generated by 3 source models: Gemini-1.5-Flash, Gemini-1.5-Pro and PaLM 2. The scoring model used is Neo-2.7 with the score function of Fast-DetectGPT. We consider five batch sizes: $k=25,50,100,250,500$. The left column displays results from the permutation test without correction, while the right column presents results of permutation test with corrected significance levels $\alpha$ for each batch test.
@@
Appendix H Experimental Results of Detecting Texts from Three Domains
@@
We also test on the dataset of Bao et al. [2023] to explore the influence of text domains on the detection result of our method. In this experiment, we only consider Scenario 1. We let $x_{t}$ be human-written text from three datasets following Mitchell et al. [2023], each chosen to represent a typical LLMs application scenario. Specifically, we incorporate news articles sourced from the XSum dataset [Narayan et al., 2018], stories from Reddit WritingPrompts dataset [Fan et al., 2018] and long-form answers written by human experts from the PubMedQA dataset [Jin et al., 2019]. Then, the capability of our algorithm is evaluated by detecting the source of texts $y_{t}$ originated from the above source models or human datasets. Source models involved in this experiment are GPT-3  [Brown, 2020], ChatGPT  [OpenAI, 2022], and GPT-4  [Achiam et al., 2023] while the scoring model is Neo-2.7 [Black et al., 2021]. The perturbation function for DetectGPT and NPR is T5-11B, and the sampling model for Fast-DetectGPT is GPT-J-6B.
@@
Figure 8(a) and Figure 8(b) present the averaged results when the two text streams are from the same domain and different domains, respectively—that is, when both sequences are sampled from or prompted using the same dataset, or different datasets. Specifically, we compute the average rejection time and false positive rate (FPR) under each significance level $\alpha$, aggregated over three LLMs (GPT-3, ChatGPT, GPT-4) and three domain configurations (three same-domain or three different-domain). A comparison of the two figures reveals that detection is faster when the prepared texts and the texts under evaluation originate from the same domain. For all score functions, our algorithm consistently controls FPRs below the target significance level $\alpha$, and the corresponding average rejection times remain below 150. Among all functions, Fast-DetectGPT is the fastest to reject $H_{0}$, with average value of $\tau$ being around 40 for same-domain texts and about 80 for different-domain texts.
@@
(a) Same domain: $x_{t}$ is sampled from XSum, $y_{t}$ is from XSum or generated by GPT-3 based on XSum.
(b) Same domain: $x_{t}$ is sampled from XSum, $y_{t}$ is from XSum or generated by ChatGPT based on XSum.
(c) Same domain: $x_{t}$ is sampled from XSum, $y_{t}$ is from XSum or generated by GPT-4 based on XSum.
(d) Same domain: $x_{t}$ is sampled from Writing, $y_{t}$ is from Writing or generated by GPT-3 based on Writing.
(e) Same domain: $x_{t}$ is sampled from Writing, $y_{t}$ is from Writing or generated by ChatGPT based on Writing.
(f) Same domain: $x_{t}$ is sampled from Writing, $y_{t}$ is from Writing or generated by GPT-4 based on Writing.
(g) Same domain: $x_{t}$ is sampled from PubMed, $y_{t}$ is from PubMed or generated by GPT-3 based on PubMed.
(h) Same domain: $x_{t}$ is sampled from PubMed, $y_{t}$ is from PubMed or generated by ChatGPT based on PubMed.
(i) Same domain: $x_{t}$ is sampled from PubMed, $y_{t}$ is from PubMed or generated by GPT-4 based on PubMed.
@@
Figure 9: Test results: mean rejection times (under $H_{1}$) and FPRs (under $H_{0}$) under each significance level $\alpha$ using 10 score functions, with texts $x_{t}$ and $y_{t}$ from the same domain. There are 3 source models: GPT-3, ChatGPT and GPT-4. Scoring model: Neo-2.7.
(a) Different domains: $x_{t}$ is sampled from XSum, $y_{t}$ is from Writing or generated by GPT-3 based on Writing.
(b) Different domains: $x_{t}$ is sampled from XSum, $y_{t}$ is from Writing or generated by ChatGPT based on Writing.
(c) Different domains: $x_{t}$ is sampled from XSum, $y_{t}$ is from Writing or generated by GPT-4 based on Writing.
(d) Different domains: $x_{t}$ is sampled from XSum, $y_{t}$ is from PubMed or generated by GPT-3 based on PubMed.
(e) Different domains: $x_{t}$ is sampled from XSum, $y_{t}$ is from PubMed or generated by ChatGPT based on PubMed.
(f) Different domains: $x_{t}$ is sampled from XSum, $y_{t}$ is from PubMed or generated by GPT-4 based on PubMed.
(g) Different domains: $x_{t}$ is sampled from Writing, $y_{t}$ is from PubMed or generated by GPT-3 based on PubMed.
(h) Different domains: $x_{t}$ is sampled from Writing, $y_{t}$ is from PubMed or generated by ChatGPT based on PubMed.
(i) Different domains: $x_{t}$ is sampled from Writing, $y_{t}$ is from PubMed or generated by GPT-4 based on PubMed.
@@
Figure 10: Test results: mean rejection times (under $H_{1}$) and FPRs (under $H_{0}$) under each significance level $\alpha$ using 10 score functions, with texts $x_{t}$ and $y_{t}$ from different domains. There are 3 source models: GPT-3, ChatGPT and GPT-4. Scoring model: Neo-2.7.
@@
Since we let $\epsilon$ equal to the actual $\Delta$ value of two sequences of human texts, which ensures FPRs of all score functions for each significance level $\alpha$ remain below $\alpha$. As we have mentioned previously, our algorithm can ensure a nonnegative supermartingale wealth under $H_{0}$ and thus control the type-I error. Besides, the relative magnitude of $\Delta-\epsilon$ and $d_{t}-\epsilon$ would influence the rejection time under $H_{1}$.
@@
Table 9: Values of $d_{t}$ with the assumption that we know the range of $|g_{t}|=\lvert\phi(x_{t})-\phi(y_{t})\rvert$ beforehand. Every two columns starting from the third column represent the $d_{t}$ values for each $t$ used under $H_{1}$ and $H_{0}$ in each test scenario.
@@
Table 10: Values of $d_{t}$ with the assumption that we know the range of $|g_{t}|=\lvert\phi(x_{t})-\phi(y_{t})\rvert$ beforehand. Every two columns starting from the third column represent the $d_{t}$ values for each $t$ used under $H_{1}$ and $H_{0}$ in each test scenario.
@@
Table 11: Values of $\Delta$ derived by using Neo-2.7 as the scoring model for the first eight score functions to score texts from the same domain. There are three source models: GPT-3, ChatGPT and GPT-4. Every two columns starting from the third column represent the $\Delta$ values under $H_{1}$ and $H_{0}$ in each test scenario.
@@
Table 12: Values of $\Delta$ derived by using Neo-2.7 as the scoring model for the first eight score functions to score texts from different domains. There are three source models: GPT-3, ChatGPT and GPT-4. Every two columns starting from the third column represent the $\Delta$ values under $H_{1}$ and $H_{0}$ in each test scenario.
@@ Appendix I Addressing Parameter Estimation Challenges and Mixed LLMs or Sources Tasks
Estimate Parameters Based on More Samples. Better parameter estimation can enhance the performance of our algorithm. For instance, in the previous experiments, we used the first $10$ samples of each sequence to estimate parameters $d_{t}$ and $\epsilon$. Here, we give an example to show that using a larger sample size for estimation could possibly yield better results. Specifically, using $20$ samples from each sequence for estimation, with test begins at the $21$-th time step, could lead to improved algorithm performance.
@@
Figure 11: Comparison of different durations in the initial stage for parameter estimation in Scenario 2. Here, text $x_{t}$ is sampled from XSum and $y_{t}$ is from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Neo-2.7. Subfigure (b) suggests that a longer duration for parameter estimation may lead to improved results. We emphasize that the test begins at $t=21$.
@@
Figure 12: (Extension to other settings) (a) Results when the sequence of texts $y_{t}$ are produced by various LLMs instead of a single one. (b) Results under the setting that the null hypothesis corresponds to the case that all the texts from the unknown source are human-written, while the alternative hypothesis $H_{1}$ corresponds to the one that not all $y_{t}$ are human-written.